# Maximum Independent Set; Hardness Parameter; Neutral Atoms; Rydberg Blockade; QuEra’s Aquila; Analog Quantum Computing; Adiabatic Schedules; Counterdiabatic Driving

### v2

@@ A.1 Nonexistence of a lower bound on the detuning
The Van der Waals interaction decays quickly with the distance between atoms, but it cannot be neglected for distances close to the blockade radius. As an example, consider the graph that consists of three atoms in a row with a distance $2a$ between its extremities, see Fig. 10(a). In this case, the MIS solution is $(1,0,1)$, which corresponds to an energy $-2\Delta_{\text{f}}+V_{0}/64$. By listing the energy of all other configurations, we find the following conditions on $\Delta_{\text{f}}$ such that |$\ket{101}$ is indeed the ground state of the final Rydberg Hamiltonian:
@@
The upper bound is of course already satisfied in Eq. (10) as it is a sufficient condition, but the value $V_{0}/64$ should be considered as a strict lower bound for any protocol; remark that this bound is mentioned in [21]. The question is then to determine if there exists, for unit disk graphs in a square lattice, a finite lower bound on the detuning to ensure the equivalence between the classical MIS solution and the ground state of the quantum Hamiltonian. Unfortunately, the answer is negative. In fact, consider the unit disk graph in Fig. 10(b), and the following two sets of vertices:
  • $S_{0}$:
    the MIS solution with $m+1$ vertices on the main line (filled circles)
  • $S_{1}$:
    the $m$ vertices located at the tip of the alternating upward and downward triangles (orange circles)
Taking into account all pairwise interactions, the corresponding energies read:
@@
where $\Sigma_{0,1}$ stands for the sums in the r.h.s. of Eq. (12). However, one can check that the difference $\Sigma_{0}-\Sigma_{1}$ is always positive and becomes arbitrarily large for increasing $m$. In particular, this difference is larger than 8 for $m\geq 9$, leading to a contradiction with Eq. (10).
In conclusion, there does not exist a value of the final detuning such that the Rydberg Hamiltonian is ensured to encode the MIS solution of all unit disk graphs.
Figure 10: Examples of unit disk graphs to study the lower bound on the final detuning. The MIS solution (filled circles) has a lower energy than the suboptimal configuration (orange circles) iff $\Delta_{\text{f}}>\Delta_{\text{LB}}$. (a) Reference graph used in Eq. (11). (b) Graph with $N=3m+1$ vertices forming $m$ alternating upward and downward triangles. For any $\Delta_{\text{f}}>0$, there exists a value of $m$ such that the suboptimal configuration is energetically favorable. Finally, two non-planar graphs with very different lower bounds: (c) $\Delta_{\text{LB}}\approx V_{0}/1017$ ($\mathcal{HP}\approx 0.42$), and (d) $\Delta_{\text{LB}}\approx V_{0}/11.49$ ($\mathcal{HP}=15$).
@@ A.2 Statistical analysis of the bounds on the detuning
Figure 11: Box plot of the lower bounds $\Delta_{\text{LB}}$ (purple) and upper bounds $\Delta_{\text{UB}}$ (yellow-green) on the detuning $\Delta_{\text{f}}$ such that the ground state of the Hamiltonian encodes the MIS solution. Calculations have been performed on 10 0001000010\,00010 000 non-isomorphic unit disk graphs with 8 to 19 vertices. All but 4 graphs satisfy $\Delta_{\text{LB}}\leq V_{0}/12$ (dashed red line) while $\Delta_{\text{UB}}$ is always strictly larger than $V_{0}/8$. It is interesting to note that, in average, easy graphs exhibit much smaller lower bounds than hard graphs (red curve), see also Fig. 10(c-d). Since the average value of $\Delta_{\text{LB}}$ quickly saturates for not-too-easy graphs, we are confident that these bounds can be effectively applied to graphs of any size.
As shown above, the final detuning has to be arbitrarily high if one wants to encode the MIS solution of any graph in the Rydberg Hamiltonian. However, this is not the case in general, as most graphs obey well-defined conditions on $\Delta_{\text{f}}$. Based on an extensive analysis of 10 0001000010\,00010 000 graphs, see Fig. 11, we find that more than 99.9% of all graphs satisfy the following bounds:
@@
Interestingly, this interval is very close to the best range of parameters found by a variational optimization, see [21, Fig. 6.1], with $\Delta_{\text{LB}}=28.91$ and $\Delta_{\text{UB}}=43.36$ MHz for QuEra’s Aquila device and $a=5$ µm. Note that a statistical analysis is also performed in [19, Sec. S2], where it is stated that about 99% of the MIS problems can be encoded in the ground state of a Rydberg Hamiltonian for well-chosen fixed experimental parameters. We strongly recommend to consider the interval in Eq. (14) when setting the final detuning of any adiabatic protocol for the MIS problem. A value which lies far outside this interval will likely lead to poor numerical or experimental results simply because the ground state of the Hamiltonian does not correspond to the MIS solution.
@@ Appendix B Graph-dependent counterdiabatic driving
For convenience, we set $\hbar=1$ and $\phi(t)=0$ in what follows, but all results can be extended to the general scenario. In this case, the Rydberg Hamiltonian reads:
@@ B.1 Approximate adiabatic gauge potential
The idea of the CD driving is to add a gauge potential $\mathcal{A}$ to the Hamiltonian $H$ such that the transitions between eigenstates are suppressed during a shorten adiabatic evolution:
@@
where $\mathcal{A}$ satisfies the following equation [27, Eq. (2)]:
@@
Neglecting the interactions $\hat{V}$, it is shown in [28] that a solution of this equation exists:
@@
This gauge potential leads to improved results compared to a simple adiabatic protocol, but one might expect even better results with a gauge potential that explicitly depends on the coefficients $V_{ij}$, hence being specific to the graph to be solved. In order to find such a generalized gauge potential, we consider the variational procedure that consists in minimizing the action [27, Eq. (8)]
$\mathcal{S}(\mathcal{A}_{\lambda})=\operatorname{Tr}\left(G^{2}(\mathcal{A}_{\lambda})\right)\quad\text{with}\quad G(\mathcal{A}_{\lambda})=\partial_{t}H+\mathrm{i}\mkern 1.0mu\ [\mathcal{A}_{\lambda},H].$ (19)
Here, we use the ansatz $\mathcal{A}_{\lambda}=-\hat{Y}\lambda(t)/2$ since the one-body Pauli matrices $\hat{Y}_{i}$ are the only operators that contribute to the action while being experimentally implementable through the phase $\phi$ of the Rabi drive. One easily checks that $\mathcal{S}$ is quadratic in $\lambda$ and that the optimum value of this parameter is:
$\lambda=2\,\frac{\operatorname{Tr}\left(\partial_{t}H\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},H]\right)}{\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},H]^{2}\right)}.$ (20)
Based on the algebraic properties of the Pauli matrices and the trace operator, this expression can be developed and simplified as follows:
@@
where the normalized traces $T_{1}$ and $T_{2}$ are given by:
$N\cdot T_{1}$ $=2^{-N}\operatorname{Tr}\left(\hat{X}\cdot[\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]\right)=\sum_{i<j}V_{ij},$ (22a)
$N\cdot T_{2}$ $=2^{-N}\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]^{2}\right)=\frac{1}{2}\sum_{i\leq j}\sum_{k\neq i,j}V_{ik}V_{kj}.$ (22b)
As expected, the graph-dependent gauge potential $\mathcal{A}_{\lambda}$ is equal to $\mathcal{A}_{V\rightarrow 0}$ in the case of vanishing interactions, with $T_{1}=T_{2}=0$ in this limit. Due to the fast-decaying interaction between distant atoms, $T_{1}$ and $T_{2}$ do not scale with $N$ as they represent average values of (nearly) local functions. For instance, $T_{1}$ is exactly half the mean degree of the underlying graph if ones considers its adjacency matrix instead of the interaction terms in Eq. (22a).
Finally, the driving amplitude and phase of the full Rydberg Hamiltonian $\widetilde{H}$ become:
@@ B.2 Practical application of the counterdiabatic driving
Hardware implementation
It follows from Eq. (23a) that $\widetilde{\Omega}$ is always larger than $\Omega$, so that it may exceed the maximum Rabi frequency $\Omega_{\max}$ of the quantum computer. In this case, we rescale the adiabatic function $\Omega$ by a factor $10<\kappa<$ 1 to satisfy the hardware constraint at any time $t$ of the evolution:
@@
Note that this is slightly different from rescaling $\widetilde{\Omega}$ directly, as the relations (21, 23) would not hold in that case.
Variational parameter
Since Eq. (21) is a generalization of the counterdiabatic term proposed in [28] through the addition of $T_{1}$ and $T_{2}$, it is interesting in an optimization context to consider a continuous interpolation between these two protocols. One possibility is to consider the interaction strength $V_{0}$ as a free parameter in the range $[0,C_{6}/a^{6}]$ when calculating the CD driving. To avoid any confusion with the fixed interactions in the Rydberg Hamiltonian, we denote by $\nu$ this free parameter. Numerically, this is equivalent to replace $T_{1}$ and $T_{2}$ in Eq. (21) by the following expressions:
@@
where $T_{1}^{(0)}$ and $T_{2}^{(0)}$ are the traces evaluated at $V_{0}=1$ in Eq. (22). Another possibility would be, e.g., to consider the two-variable parametrization $T_{i}=\nu_{i}\,T_{i}^{(0)}$, with $\nu_{i}\in[0,V_{0}^{i}]$ and $i\in\{1,2\}$.
@@ Appendix C Graph datasets and numerical implementation
In the following sections, we first describe two sets of unit disk graphs $\mathcal{G}=(V,E)$ that are used in this study; the data are available on GitHub [40]. In all examples, we consider an underlying square lattice where the distance between nearest-neighbors is $a=1$ and where any pair of vertices in $V$ is connected by an edge in $E$ if the corresponding distance is smaller than or equal to $\sqrt{2}$. Then, we detail the numerical method that is used to simulate the evolution of the quantum state under the action of the Rydberg Hamiltonian.
@@ C.1 Graph datasets
From 10 0001000010\,00010 000 non-isomorphic graphs randomly generated, we choose 500 graphs such that the correlation between their hardness parameter and their order is as small as possible, see Fig. 12. More precisely, we apply the following rules to determine which graphs belong to the dataset:
  • there are exactly 50 graphs for each order from 8 to 17
  • the hardness parameters are distributed as evenly as possible in a log-scale from 0.375 to 12; in particular, no two graphs share the same hardness parameter, and exactly 100 graphs are contained in each bin $[3\cdot 2^{i-4},3\cdot 2^{i-3})$ with $i=1,\ldots,5$
  • the distribution of the graph orders is as uniform as possible within each such bin
Figure 12: Visualization of the hardness parameters $\mathcal{HP}$ and the orders of the 500 small graphs. Due to the combinatorial definition of the hardness parameter, graphs with fewer vertices (down purple triangles) cannot generate very hard MIS instances. This explains why graphs of order 13 and more (up yellow-green triangles) are overrepresented in the range $\mathcal{HP}\in[6,12]$.
@@
For completeness, we display in Fig. 13 the 11 “toy graphs” defined in [29, Appendix C]:
  • the underlying square lattice has $4\times 3$ sites
  • the order of the graphs is 9 or 10
  • there is only one MIS solution for each graph
Figure 13: The 11 “toy graphs” sorted by the number of vertices (9 or 10) and by hardness parameter (between 2 and 3). The unique MIS is highlighted with dark blue circles. To refer to a specific graph, we generate the three-nibble identifiers as follows: the $i$-th nibble (four-bit aggregation) is the hexadecimal symbol that represents the absence or presence of a vertex in the 4 sites of the $i$-th row (from top to bottom).
@@ C.2 Numerical implementation
Several software development kits (SDK) exist to simulate the Rydberg evolution: Bloqade [43] and Amazon Braket [32] for QuEra’s hardware, or Pulser [33] based on QuTiP for Pasqal’s technology, just to name a few. Nevertheless, we developed our own routines to get a full control over all parameters and, more importantly, to fit exactly our needs. In fact, optimizing hundreds of graphs for several protocols requires intense (classical) computing power, so that we have to find the perfect balance between accuracy and speed. For example, for each graph of the datasets we precompute all time- and protocol-independent terms in the Hamiltonian, namely the $\hat{X}$, $\hat{Y}$, $\hat{n}$ and $\hat{V}$ operators, see Eq. (15). We also use an explicit Runge-Kutta method of order 2 [44] to integrate the time-dependent Schrödinger equation, which turns out to be sufficiently accurate and much faster than the default order 8 method (DOP853 in Python or Vern8 in Julia). The initial state of this ordinary differential equation is |$\ket{0}^{\otimes N}$, which corresponds to the unique ground state of the Rydberg Hamiltonian with $\Omega=0$ and $\Delta<0$.
Blockade subspace

### v3

@@ A.1 Nonexistence of a lower bound on the detuning
The van der Waals interaction decays quickly with the distance between atoms, but it cannot be ignored for distances close to the blockade radius. As an example, consider the graph that consists of three atoms in a row with distance $2a$ between its extremities; see Fig. 11(a). In this case, the MIS solution is $(1,0,1)$, which corresponds to an energy of $-2\Delta_{\text{f}}+V_{0}/64$. By listing the energy of all other configurations, we find the following conditions on $\Delta_{\text{f}}$ such that |$\ket{101}$ is indeed the ground state of the final Rydberg Hamiltonian:
@@
The upper bound is, of course, already satisfied in Eq. (10) as it is a sufficient condition, but the value $V_{0}/64$ should be considered as a strict lower bound for any protocol; note that this bound is mentioned in Ref. [25]. The question is then to determine if there exists, for unit disk graphs in a square lattice, a finite lower bound on the detuning to ensure the equivalence between the classical MIS solution and the ground state of the quantum Hamiltonian. Unfortunately, the answer is negative. Consider the unit disk graph in Fig. 11(b), and the following two sets of vertices: $S_{0}$, the MIS solution with $m+1$ vertices on the main line (filled circles); $S_{1}$, the $m$ vertices located at the tip of the alternating upward and downward triangles (orange circles). If we take into account all pairwise interactions, the corresponding energies read
@@
where $\Sigma_{0}$ and $\Sigma_{1}$ stand for the sums on the right-hand sides of Eqs. (12a) and (12b). However, one can check that the difference $\Sigma_{0}-\Sigma_{1}$ is always positive and becomes arbitrarily large for increasing $m$. In particular, this difference is larger than 8 for $m\geq 9$, leading to a contradiction with Eq. (10).
In conclusion, there does not exist a value of the final detuning such that it is ensured that the Rydberg Hamiltonian encodes the MIS solution of all unit disk graphs.
Figure 11: Examples of unit disk graphs to study the lower bound on the final detuning. The MIS solution (filled circles) has a lower energy than the suboptimal configuration (orange circles) if and only if $\Delta_{\text{f}}>\Delta_{\text{LB}}$. (a) Reference graph used in Eq. (11). (b) Graph with $N=3m+1$ vertices forming $m$ alternating upward and downward triangles. For any $\Delta_{\text{f}}>0$, there exists a value of $m$ such that the suboptimal configuration is energetically favorable. (c), (d) Two non-planar graphs with very different lower bounds: (c) $\Delta_{\text{LB}}\approx V_{0}/1017$ ($\mathbb{H}\approx 0.42$) and (d) $\Delta_{\text{LB}}\approx V_{0}/11.49$ ($\mathbb{H}=15$).
@@ A.2 Statistical analysis of the bounds on the detuning
Figure 12: Box plot of the lower bounds $\Delta_{\text{LB}}$ (purple) and upper bounds $\Delta_{\text{UB}}$ (yellow-green) on the detuning $\Delta_{\text{f}}$ such that the ground state of the Hamiltonian encodes the MIS solution. Calculations are performed on 10 0001000010\,00010 000 nonisomorphic unit disk graphs with 8–19 vertices. All but four graphs satisfy $\Delta_{\text{LB}}\leq V_{0}/12$ (dashed red line) while $\Delta_{\text{UB}}$ is always strictly larger than $V_{0}/8$. It is interesting to note that, on average, easy graphs exhibit much smaller lower bounds than hard graphs (red curve); see also Fig. 11(c-d). Since the average value of $\Delta_{\text{LB}}$ quickly saturates for not-too-easy graphs, we are confident that these bounds can be effectively applied to graphs of any size.
As shown above, the final detuning has to be arbitrarily high if one wants to encode the MIS solution of any graph in the Rydberg Hamiltonian. However, this is not the case in general, as most graphs obey well-defined conditions on $\Delta_{\text{f}}$. From an extensive analysis of 10 0001000010\,00010 000 graphs—see Fig. 12—we find that more than 99.9% of all graphs satisfy the following bounds:
@@
Interestingly, this interval is very close to the best range of parameters found by a variational optimization (see Fig. 6.1 in Ref. [25]) with $\Delta_{\text{LB}}=28.91$ MHz and $\Delta_{\text{UB}}=43.36$ MHz for QuEra’s Aquila device and $a=5$ µm. Note that a statistical analysis is also performed in Ref. [19] (see Sec. S2 therein), where it is stated that about 99% of the MIS problems can be encoded in the ground state of a Rydberg Hamiltonian for well-chosen fixed parameters of the quantum hardware. We strongly recommend considering the interval in Eq. (14) when one is setting the final detuning of any adiabatic protocol for the MIS problem. A value that lies far outside this interval will likely lead to poor results simply because the ground state of the Hamiltonian does not correspond to the MIS solution.
@@
For convenience, we set $\hbar=1$ and $\phi(t)=0$ in what follows, but all results can be extended to the general scenario. In this case, the Rydberg Hamiltonian reads
@@ B.1 Approximate adiabatic gauge potential
The idea of the CD driving is to add a gauge potential $\mathcal{A}$ to the Hamiltonian $H$ such that the transitions between eigenstates are suppressed during a shortened adiabatic evolution:
@@
where $\mathcal{A}$ satisfies the equation [see Eq. (2) in Ref. [31]]:
@@
If the interactions $\hat{V}$ are ignored, it is shown in Ref. [32] that a solution of this equation exists:
@@
This gauge potential leads to improved results compared with a simple adiabatic protocol, but one might expect even better results with a gauge potential that explicitly depends on the coefficients $V_{ij}$, hence being specific to the graph to be solved. To find such a generalized gauge potential, we consider the variational procedure that consists in minimizing the action [see Eq. (8) in Ref. [31]]
$\mathcal{S}(\mathcal{A}_{\lambda})=\operatorname{Tr}\left(G^{2}(\mathcal{A}_{\lambda})\right)\text{, with }G(\mathcal{A}_{\lambda})=\partial_{t}H+\mathrm{i}\mkern 1.0mu[\mathcal{A}_{\lambda},H].$ (19)
Here we use the ansatz $\mathcal{A}_{\lambda}=-\hat{Y}\lambda(t)/2$ since the one-body Pauli matrices $\hat{Y}_{i}$ are the only operators that contribute to the action while being implementable in the quantum hardware through the phase $\phi$ of the Rabi drive. One can easily check that $\mathcal{S}$ is quadratic in $\lambda$ and that the optimum value of this parameter is given by
$\lambda=2\,\frac{\operatorname{Tr}\left(\partial_{t}H\,[\mathrm{i}\mkern 1.0mu\hat{Y},H]\right)}{\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},H]^{2}\right)}.$ (20)
On the basis of the algebraic properties of the Pauli matrices and the trace operator, this expression can be developed and simplified as follows:
@@
where the normalized traces $T_{1}$ and $T_{2}$ are given by
$N\,T_{1}$ $=2^{-N}\operatorname{Tr}\left(\hat{X}\,[\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]\right)=\sum_{i<j}V_{ij},$ (22a)
$N\,T_{2}$ $=2^{-N}\operatorname{Tr}\left([\mathrm{i}\mkern 1.0mu\hat{Y},\hat{V}]^{2}\right)=\frac{1}{2}\sum_{i\leq j}\sum_{k\neq i,j}V_{ik}V_{kj}.$ (22b)
As expected, the graph-dependent gauge potential $\mathcal{A}_{\lambda}$ is equal to $\mathcal{A}_{V\rightarrow 0}$ in the case of vanishing interactions, with $T_{1}=T_{2}=0$ in this limit. Because of the fast-decaying interaction between distant atoms, $T_{1}$ and $T_{2}$ do not scale with $N$ as they represent average values of (nearly) local functions. For instance, $T_{1}$ is exactly half the mean degree of the underlying graph if one considers its adjacency matrix instead of the interaction terms in Eq. (22a).
Finally, the driving amplitude and phase of the full Rydberg Hamiltonian $\widetilde{H}$ become
@@ B.2 Practical application of the counterdiabatic driving
B.2.1 Hardware implementation
It follows from Eq. (23a) that $\widetilde{\Omega}$ is always larger than $\Omega$, so it may exceed the maximum Rabi frequency $\Omega_{\max}$ of the quantum computer. In this case, we rescale the adiabatic function $\Omega$ by a factor $10<\kappa<$ 1 to satisfy the hardware constraint at any time $t$ of the evolution:
@@ B.2.1 Hardware implementation
Note that this is slightly different from rescaling $\widetilde{\Omega}$ directly, as the relations (21), (23a), and (23b) would not hold in that case.
B.2.2 Variational parameter
Since Eq. (21) is a generalization of the counterdiabatic term proposed in Ref. [32] through the addition of $T_{1}$ and $T_{2}$, it is interesting in an optimization context to consider a continuous interpolation between these two protocols. One possibility is to consider the interaction strength $V_{0}$ as a free parameter in the range $[0,C_{6}/a^{6}]$ when one is calculating the CD driving. To avoid any confusion with the fixed interactions in the Rydberg Hamiltonian, we denote by $\nu$ this free parameter. Numerically, this is equivalent to replacing $T_{1}$ and $T_{2}$ in Eq. (21) by the following expressions:
@@ B.2.2 Variational parameter
where $T_{1}^{(0)}$ and $T_{2}^{(0)}$ are the traces evaluated at $V_{0}=1$ in Eqs. (22a) and (22b). Another possibility would be to consider the two-variable parametrization $T_{i}=\nu_{i}\,T_{i}^{(0)}$, with $\nu_{i}\in[0,V_{0}^{i}]$ and $i\in\{1,2\}$.
@@
In the following sections, we first describe two sets of unit disk graphs $\mathcal{G}=(V,E)$ that are used in this study; the data are available on GitHub [48]. In all examples, we consider an underlying square lattice where the distance $a$ between nearest-neighbors is $1$ and where any pair of vertices in $V$ is connected by an edge in $E$ if the corresponding distance is smaller than or equal to $\sqrt{2}$. Then we detail the numerical method that is used to simulate the evolution of the quantum state under the action of the Rydberg Hamiltonian.
@@ C.1 Graph datasets
From 10 0001000010\,00010 000 nonisomorphic graphs randomly generated, we choose 500 graphs such that the correlation between their hardness parameter and their order is as small as possible; see Fig. 13. More precisely, we apply the following rules to determine which graphs belong to the dataset:
  • There are exactly 50 graphs for each order from 8 to 17.
  • The hardness parameters are distributed as evenly as possible on a logarithmic scale from 0.375 to 12; in particular, no two graphs share the same hardness parameter, and exactly 100 graphs are contained in each bin $[3\times 2^{i-4},3\times 2^{i-3})$ with $i=1,\ldots,5$.
  • The distribution of the graph orders is as uniform as possible within each such bin.
Figure 13: Visualization of the hardness parameters $\mathbb{H}$ and the orders of the 500 small graphs. Because of the combinatorial definition of the hardness parameter, graphs with few vertices (6–12 purple triangles) cannot generate very hard MIS instances. This explains why larger graphs (13–17 vertices, yellow-green triangles) are overrepresented in the range $\mathbb{H}\in[6,12]$.
@@
For completeness, we display in Fig. 14 the 11 toy graphs defined in Appendix C in  [33]:
  • The underlying square lattice has $4\times 3$ sites.
  • The order of the graphs is 9 or 10.
  • There is only one MIS solution for each graph.
Figure 14: The 11 toy graphs sorted by the number of vertices (nine or ten) and by hardness parameter (between 2 and 3). The unique MIS is highlighted with dark-blue circles. To refer to a specific graph, we generate the three-nibble identifiers as follows: the $i$th nibble (4-bit aggregation) is the hexadecimal symbol that represents the absence or presence of a vertex in the four sites of the $i$th row (from top to bottom).
@@ C.2 Numerical implementation
Several software development kits exist to simulate the Rydberg evolution: for example, BLOQADE [51] and AMAZON BRAKET [40] for QuEra’s hardware, or PULSER [41] based on QuTiP for Pasqal’s technology. Nevertheless, we developed our own routines to get full control over all parameters and, more importantly, to fit exactly our needs. Optimization of hundreds of graphs for several protocols requires intense (classical) computing power, so we have to find the perfect balance between accuracy and speed. For example, for each graph of the datasets we precompute all time- and protocol-independent terms in the Hamiltonian, namely, the $\hat{X}$, $\hat{Y}$, $\hat{n}$, and $\hat{V}$ operators; see Eq. (15). We also use an explicit Runge-Kutta method of order 2 [52] to integrate the time-dependent Schrödinger equation, which turns out to be sufficiently accurate and much faster than the default order 8 method (DOP853 in PYTHON or VERN8 in JULIA). The initial state of this ordinary differential equation is |$\ket{0}^{\otimes N}$, which corresponds to the unique ground state of the Rydberg Hamiltonian with $\Omega=0$ and $\Delta<0$.
C.2.1 Blockade subspace